# R-LoRA: Random Initialization of Multi-Head LoRA for Multi-Task Learning

### v1

@@ 4.2 Multi-Head Random Initialization
Metrics Base LoRA LoRAHub* LoRA MoE* HydraLoRA R-LoRA
7B 31.6 37.2 39.7 40.3 41.8 42.3
13B 38.4 40.9 41.9 43.7 44.7 45.4
A/B for training - 1/1 48/48 48/48 1/10 1/10
A/B for inference - 1/1 20/20 48/48 1/10 1/10
% Param - 0.062 1.240 2.976 0.341 0.341
Table 3: Comparison of different training schemes on multi task. * indicates results from Tian et al. (2024)
Figure 6: Gradient norms of HydraLoRA and R-LoRA.
5 Experiment
In this section, we validate the superiority of R-LoRA across various models and settings. First, we followed the settings of Tian et al. (2024) and conducted experiments on the LLaMA-2 model Touvron et al. (2023), evaluating both single-task and multi-task scenarios. Subsequently, we tested the performance of R-LoRA under different multi-task settings on the new Qwen2.5 Qwen Team (2024). The model sizes range from 0.5B to 13B. Through an extensive ablation study, we demonstrate the effectiveness of the multi-head randomization in R-LoRA.
5.1 Experiment Setting
Model: In the single-task setting, we use LLaMA2-7B, while in the multi-task setting, we additionally incorporated LLaMA2-13B. In the ablation study, we use Qwen2.5-0.5B and Qwen2.5-3B models.
Dataset & Benchmarks:
Single-task:
  • General: We fine-tune the model using the general instruction tuning dataset Databricks-Dolly-15k Conover et al. (2023), which focuses on generic language capabilities. The performance is then evaluated using the MMLU benchmark Hendrycks et al. (2020).
  • Medical: We fine-tune the model using GenMedGPT and the Clinic-10k dataset from ChatDoctor Li et al. (2023b), targeting medical applications. The model’s performance on medical tasks is assessed using MMLU.
  • Law: Fine-tuning is conducted using two legal instruction datasets, Lawyer-Instruct Alignment-Lab-AI (2023) and US-Terms Chalkidis et al. (2023), and the model is evaluated on legal tasks using MMLU.
  • Code: We fine-tune the model using the CodeAlpaca Chaudhary (2023) for code generation tasks, and the evaluation is conducted using the HumanEval Chen et al. (2021).
  • Math: The model is fine-tuned on the training set of GSM8K Cobbe et al. (2021) to enhance its mathematical reasoning capabilities and is evaluated on the corresponding test set
Multi-task: We fine-tune the model using a subset of the Flanv2 dataset Brown (2020) that includes tasks from both Natural Language Understanding (NLU) and Natural Language Generation (NLG), grouped into 10 distinct task clusters. The model’s performance is evaluated using the Big-Bench Hard (BBH) benchmark. More Details of the datasets will be provided in the appendix A.
Baseline: First, we compare R-LoRA against various PEFT methods on single datasets: 1) Full fine-tuning; 2) Prompt Tuning Lester et al. (2021); 3) P-Tuning Liu et al. (2024c); 4) Prefix Tuning Li and Liang (2021); 5) $IA^{3}$ Liu et al. (2022); 6) AdaLoRA Zhang et al. (2023); 7) HydraLoRA Tian et al. (2024). Second, we extend the comparison by evaluating R-LoRA against other weighted averaging methods across multiple datasets: 1) Lorahub Huang et al. (2023), which utilizes black-box optimization to learn weights for 20 randomly selected LoRAs for new tasks, applying weighted averaging without the need for gradient calculations; 2) LoRA MoE Liu et al. (2024a), which combines lightweight experts (LoRA) with a Mixture of Experts (MoE) architecture for high efficiency, enabling generalization to new tasks without prior knowledge; 3) HydraLoRA Tian et al. (2024), which employs Multi-Head structure in conjunction with MoE to achieve a balance between parameter efficiency and training effectiveness.
5.2 Performance
5.2.1 Performance of R-LoRA on Single Task
As shown in Table 2, in the single-task setting, where the knowledge and text format of the data are relatively homogeneous, multi-head randomization does not yield significant performance improvements. Nevertheless, R-LoRA achieves performance on par with HydraLoRA, demonstrating that the multi-head randomization mechanism preserves learning effectiveness while maintaining stability for single-task scenarios. This highlights R-LoRA’s robustness and adaptability, even in settings where its full potential may not be fully utilized.
5.2.2 Performance of R-LoRA on Multi-Tasks
The evaluation across diverse tasks, as shown in Table 3, demonstrates that R-LoRA, building upon the foundation of HydraLoRA, consistently outperforms all other schemes. By introducing multi-head dropout and random initialization for the head matrices, R-LoRA further enhances the model’s stability and adaptability. The performance gains of R-LoRA, rooted in these multi-head randomization techniques, surpass those of both conventional PEFT methodologies and HydraLoRA.
@@ 5.2.2 Performance of R-LoRA on Multi-Tasks
+MT 91.20 80.80 77.50 66.20 49.40 73.02
+Init 91.40 81.20 77.10 66.10 49.10 72.98
R-LoRA 91.74 81.50 77.60 67.80 49.30 73.59
Table 4: Results of Ablation Studies on Qwen2.5-0.5B with Different Schemes Across Various Tasks. The table compares the performance of HydraLoRA, +MT (HydraLoRA with Multi-Head Dropout), +Init (Multi-Head Random Initialization), and R-LoRA across five tasks.
Schemes Task1 2 3 4 5 6 7 8 Avg
Qwen2.5-0.5B
HydraLoRA 87.5 77 73.5 63.5 52.5 52 54.5 52.5 64.13
+MT 88 78.5 74.5 66.5 54.75 53.5 58.5 52.5 66.22
+Init 87.5 79.75 75.75 65.75 53 58 56 55 66.31
R-LoRA 89 82.5 76.5 66.5 55 57 55.5 56.5 67.31
Qwen2.5-3B
HydraLoRA 95.5 84 83.5 83.5 71.25 72 83.5 88 82.66
+MT 96 83.5 83.5 84.5 71.75 72 83.5 88 82.84
+Init 96.5 84 83.5 84.75 71.5 73 85 88.5 83.34
R-LoRA 96.5 83.5 85 86.5 72.75 73.5 86 89 84.09
Table 5: Results of Ablation Studies on Qwen2.5-0.5B and Qwen2.5-3B Models with Different Schemes Across Various Tasks. The table compares the performance of HydraLoRA, +MT (HydraLoRA with Multi-Head Dropout), +Init (Multi-Head Random Initialization), and R-LoRA across eight tasks.
5.3 Parameter Analysis
Reserach Question2: Does multi-head randomization effectively enhance the acquisition of diverse knowledge across the head matrices?
In this section, we analyze the parameter differences among the head matrices in R-LoRA. The methodology and experimental setup align with those described in Section 3. As shown in Figure 5, the parameter similarity between head matrices in R-LoRA is reduced to below 70%. This significant decrease indicates that multi-head randomization effectively enhances the model’s capacity to learn task-specific knowledge, thereby mitigating redundant learning and increasing the diversity of acquired knowledge across tasks. T-SNE analysis will be shown in appendix C
5.4 Training Process
Reserach Question3: Does multi-head randomization impact the stability of the training process?
As illustrated in Figure 6, R-LoRA benefits from multi-head randomization, exhibiting significantly larger gradient norms in the early stages of training compared to HydraLoRA. This drives the head matrices to converge to distinct regions, enhancing the model’s ability to capture diverse representations and improving overall performance. Furthermore, R-LoRA demonstrates greater training stability than HydraLoRA, as evidenced by its more stable gradient norms throughout the training process. This stability enables the model to effectively acquire diverse knowledge without compromising training efficiency.
5.5 Ablation Study
In this section, we empirically validate the effectiveness of R-LoRA’s multi-head randomization components through extensive experiments. Ablation studies were conducted on two models, Qwen2.5-0.5B and Qwen2.5-3B, under two task settings: 5-task and 8-task configurations. For the 5-task setting, models were fine-tuned on 5 datasets spanning two categories (NLU and commonsense reasoning). For the 8-task setting, models were trained on 11 datasets across 8 categories, with all models evaluated on their respective test sets. The experimental results are presented in Table 4 and Table 5. Details of the datasets will be provided in the appendix A.3.
Experimental results demonstrate that the two key components of multi-head randomization—random initialization and dropout—are pivotal for enhancing the model’s adaptability across tasks. Random initialization assigns unique weights to each head matrix, enabling the capture of task-specific patterns and reducing the risk of head convergence. Dropout diversifies inputs to the head matrices, fostering distinct learning pathways and reducing redundancy. Together, these components improve task-specific feature capture while ensuring robustness in multi-task learning.
6 Conclusion
In this work, we first investigated the multi-head structure of LoRA and analyzed the parameters of the head matrices, revealing that they remain highly similar. To address this, we proposed R-LoRA, which introduces multi-head randomization—a simple yet effective approach—to enable the model to learn knowledge from different tasks, thereby enhancing its performance in multi-task scenarios. This method not only improves the model’s generalization capabilities but also supports its adaptability across diverse tasks. Extensive experiments have validated the superiority of R-LoRA. Parameter analysis demonstrates that multi-head randomization effectively differentiates the head matrices, enabling them to learn knowledge from distinct tasks. This capability significantly enhances the model’s performance in multi-task scenarios, confirming the effectiveness of the proposed approach.
7 Limitation
Despite the promising results of R-LoRA, several limitations should be acknowledged. While empirical evidence supports the effectiveness of multi-head randomization, a rigorous theoretical analysis of its underlying mechanisms remains absent. Additionally, multi-head random initialization does not ensure consistency with the pre-trained model’s outputs, potentially introducing random disturbances. Future work could explore data-driven initialization as a promising approach to enhance the learning of task-specific knowledge by the head matrices, a direction we intend to pursue further.
Appendix A Datasets
A.1 Single-task
@@
A.2 Multi-task
For complex mixed multi-task/domain, we select a portion of the Flanv2 datasets covering Natural Language Understanding (NLU) and Natural Language Generation (NLG), which can be grouped into 10 distinct task clusters. Then we evaluate it with the Big-Bench Hard (BBH) benchmark.
@@ A.2 Multi-task
A.3 Ablation Study
Due to limited computational resources, we selected a subset of the dataset for training and testing. Five tasks:
@@ A.3 Ablation Study
  • Task 7: Coreference Resolution (SiQA)
@@
Appendix B Baselines
@@
Appendix C More Results
The T-SNE analysis of R-LoRA has been shown in Figure 789
Figure 7: T-SNE analysis of head matrices(down_proj) in R-LoRA.
Figure 8: T-SNE analysis of head matrices(up_proj) in R-LoRA.
Figure 9: T-SNE analysis of head matrices(gate_proj) in R-LoRA.
Figure 10: Training loss curves of HydraLoRA and R-LoRA. The loss of M-LoRA remains lower throughout the entire training process..

### v2

@@ A.4 Datasets in Single-task
Table 7: Results of Ablation Studies on Qwen2.5-0.5B with Different Schemes Across Various Tasks. The table compares R-LoRA with its ablated versions (without Multi-Head Dropout/MD and without Multi-Head Random Initialization/MI) against HydraLoRA across five tasks.
Appendix B Datasets
B.1 Motivation
In the section of Motivation, We fine-tune Qwen2.5-3B on five tasks: Paraphrase Detection (QQP), Natural Language Inference (QNLI) Wang (2018), Commonsense Reasoning (SIQA) Sap et al. (2019), Physical Commonsense Reasoning (PIQA) Bisk et al. (2020), and Math (GSM8K) Cobbe et al. (2021)
B.2 Setting 1
@@
    Commonsense Reasoning: Hellaswag
  6. 6.
    Open-Book Question Answering: OBQA
  7. 7.
    Closed-Book Question Answering: ARC
  8. 8.
    Mathematical Reasoning: GSM8K
B.3 Setting 2
Following Tian et al. (2024), for complex mixed multi-task/domain, we select a portion of the Flanv2 datasets covering Natural Language Understanding (NLU) and Natural Language Generation (NLG), which can be grouped into 10 distinct task clusters. Then we evaluate it with the Big-Bench Hard (BBH) benchmark.
@@ B.3 Setting 2
B.4 Ablation Study
Due to limited computational resources, we selected a subset of the dataset for training and testing. Five tasks for Smaller model Qwen2.5-0.5B in Appendix A.3:
@@ B.4 Ablation Study
  • Task 7: Coreference Resolution (Winogrande)
@@
Appendix C Implementation Details
The hyperparameters used for training are as follows: a learning rate of 0.0002, "lora_alpha"=32, and trainable LoRA components including "gate_proj", "down_proj", and "up_proj". A dropout rate of 0.2 was applied to the LoRA, with a warmup ratio of 0.03. Mixed-precision training was enabled using bfloat16, and the learning rate scheduler was set to cosine annealing. The model was trained for 1 epoch on NVIDIA 4090 GPU. In Setting 1, the rank of LoRA was set to 10 for LoRA*2 to match the total number of trainable parameters in R-LoRA, while the rank for others was set to 4.
Figure 7: Training loss curves of HydraLoRA and R-LoRA. The loss of R-LoRA remains lower throughout the entire training process.
Appendix D Baselines